OpenStudy (anonymous):

a boats crew rowed 20 miles downstream , with the current , in 2 hours. the return trip upstream , against the current , covered the same distance , but took 5 hours. the crew rowing rate in still water is ? miles per hour the rate of the current is ? miles per hour

OpenStudy (anonymous):

\(Rate.Speed=\frac{spead1 + spead2}{2}\) speed1 = 20/2=10 mile per hour speed2= 20/5=4 mile per hour \(Rate.Speed=\frac{10 +4}{2}= 7 ...MilePerHour\) you are welcome

OpenStudy (anonymous):

The Rate of the current : let B : Boat speed r : river speed 10 = B +r ..............1 4=B-r.....................2 solve 1 and 2 to get B = 7 ... which is obtained before r = 3 mile per hour which is the current speed